OpenStudy (anonymous):

At an archery competition, Jamie and Marshall's team tied Shawna and Wyatt's team for the top team score. Team scores are determined by adding the two team member's individual scores. Marshall scored 12 less than Jamie, Shawna scored twice as much as Marshall, and Wyatt scored 4 less than Shawna. What was each person's individual score?

OpenStudy (zzr0ck3r):

we use j, m, s, w so j+m = s+w m = j-12 s = 2m w = s-4 m+12 = j m = s/2 j+m=s+w m+12+(s/2) = s+(s-4) (s/2)+12+(s/2) = 2s-4 s+12 = 2s - 4 s = 16 so w = 12 m = 8 and j = 20 j+m = s+w 20+8 = 16+12 28=28

OpenStudy (zzr0ck3r):

make sense @ksapphire12 ?

OpenStudy (anonymous):

Yes, thank you

OpenStudy (zzr0ck3r):

np